OpenStudy (anonymous):

sin2x = cosx 1. give general solutions?? 2. solve for x [-360;360]

OpenStudy (amistre64):

2sinx = 1 ... if you insist on dividng

OpenStudy (waleed_imtiaz):

sinx=1/2 yes.... u are right.... amistre

OpenStudy (amistre64):

and assuming that sin(2x) as opposed to the copy/paste version of sin^2(x)

OpenStudy (anonymous):

plzzzzz explain in detail?

OpenStudy (amistre64):

the detail is sin(2x) has an identity in terms of cos(x); so use it ....

OpenStudy (amistre64):

if its sin^2 (x), then use 1-cos^2 (x) and solve a quadratic

OpenStudy (anonymous):

could u plz saw me steps of doing it?

OpenStudy (amistre64):

you mean write out a detailed script that covers everything that you can then just copy and paste? no

OpenStudy (anonymous):

i am realy struggling with it????

OpenStudy (amistre64):

please ask questions about what you do not understand instead of asking for some full solutions how have you started it, what have you tried in working with it .... what is your effort in this?

OpenStudy (anonymous):

x = 150 +360k is my answer???

OpenStudy (amistre64):

* and .. the material that you posted is vague and ambiguous due to people breaking form and typing exponents without the "^"

OpenStudy (amistre64):

that does not show us any of your workings ... that seems more like you have some options and are guessing at them.

OpenStudy (amistre64):

but, it does seem about right ... sin(30) = 1/2; and sin(150) = sin(30) but that is not the only solution; since cos(x) = 0 when 2sin(x)cos(x) = 0

OpenStudy (anonymous):

cosx(2sin-1)=0

OpenStudy (amistre64):

when does cos(x) = 0?

OpenStudy (anonymous):

x =90 +180k

OpenStudy (amistre64):

good, now theres just one more solution set to worry about i believe

OpenStudy (amistre64):

notice when i said that sin(150) = sin(30) ?

OpenStudy (anonymous):

150+360

OpenStudy (amistre64):

yeah, but that 150 + 36ok only covers every 150 .... we are missing the 30s

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

does 150 = 30?

OpenStudy (anonymous):

explain this?

OpenStudy (amistre64):

.... does sin(150) = sin(30) ???

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

are they the same angle? or just the same sin values?

OpenStudy (anonymous):

yes 1/2

OpenStudy (amistre64):

that doesnt answer my question ....

OpenStudy (amistre64):

is the angle 150 equal to the angle 30?